### v1

@@ 5.2 Evaluation Results on LasHer
Table 1: The overall performance on LasHer. The red, blue, and green fonts indicate the first, second, and third places of each evaluation metric, respectively.
@@
M3PT Prompt fine-tune 67.3 63.9 54.2 46.1fps
ViPT-deep[17] Prompt fine-tune 64.8 61.3 52.2 -
ViPT-shaw[17] Prompt fine-tune 59.0 55.7 47.6 -
ProTrack[16] Prompt fine-tune 53.8 49.8 42.0 -
ECMD[42] Full fine-tune 59.0 54.6 46.4 30.0fps
MFNet[14] Full fine-tune 59.7 55.4 46.7 18.0fps
HMFT[15] Full fine-tune 43.6 38.1 31.3 25.1fps
mfDiMP[5] Full fine-tune 44.7 39.5 34.3 22.0fps
SGT++[43] Full fine-tune 36.5 30.6 25.1 1.2fps
PLASSO[45] Other 54.7 - 47.2 -
APFNet[13] Other 50.0 43.9 36.2 1.9fps
MANet++[8] Other 46.7 40.8 31.7 16.1fps
CAT[6] Other 45.0 39.5 31.4 -
FANet[46] Other 44.1 38.4 30.9 12.0fps
DAPNet[47] Other 43.1 38.3 31.4 -
We evaluate the performance of our proposed method on LasHer, and compare it with 14 state-of-the-art RGB-T tracking methods. These methods fall into three categories: the first category consists of methods that use pre-trained RGB tracking models for prompt fine-tuning, containing M3PT, ViPT-deep[17], ViPT-shaw[17], and ProTrack[16]; the second category comprises methods that use pre-trained RGB tracking models for full fine-tuning, containing ECMD[42], MFNet[14], HMFT[15], mfDiMP[5], and SGT++[43]; the third category includes methods that do not use pre-trained upstream parameters, containing APFNet[13], MANet++[8], FANet[46], PLASSO[45], CAT[6], and DAPNet[47]. The evaluation metrics used by LasHer are prediction accuracy (PR), normalized prediction accuracy (NPR), and success rate (SR).
Table 1 shows the evaluation results of all methods on LasHer. Our proposed M3PT method achieves PR, NPR, and SR scores of 67.3, 63.9, and 54.2 respectively, which are 2.5, 2.6, and 2.0 percentage points higher than the current best prompt fine-tuning method ViPT-deep respectively. Our method’s performance ranks first among all existing prompt fine-tuning methods. This demonstrates the superiority of our proposed method over the previous prompt methods. Moreover, our method also outperforms most other state-of-the-art methods, even including multiple fine-tune state-of-the-art methods. This indicates that our method further unleashes the huge potential of prompt fine-tuning in RGB-T tracking tasks through a reliable framework and multiple flexible prompt strategies. Furthermore, our method’s inference speed reaches 46.1fps, proving the efficiency of our proposed middle fusion tracking framework. In addition, we also use the LasHer toolkit to plot the PR, NPR, and SR evaluation curves of our method and 10 open-source state-of-the-art methods, as shown in Figure 6. Our method shows great competitiveness under all thresholds of each evaluation curve.
We notice that LasHer contains 19 classic challenge attributes and has been finely annotated, which can cover most complex scenarios in the real world. To evaluate the performance of our proposed method in these scenarios and further verify whether it can fully exploit the advantages of RGB-T tracking, we use the LasHer toolkit to plot per-attribute PR evaluation curves of our method and 10 other open-source methods under these 19 challenge attributes. The evaluation results are shown in Figure 7. Our method shows leading performance that far exceeds other similar methods on 18 challenge attributes (TO, TC, OV, SV, PO, NO, SA, MB, LI, HO, ARC, LR, CM, FM, HI, BC, IV, DEF), which include not only the common challenge attributes of visual tracking (TO, OV, SV, PO, NO, SA, MB, ARC, LR, CM, FM, BC, DEF), but also the challenge attributes that cause serious degradation of the reliability of the visible modality (LI, HI, AIV) and thermal infrared modality (TC). This proves that our method can indeed fully explore and utilize the key modes of modality sharing and modality independence, achieving robust tracking in various challenge scenarios.
@@ 5.3 Evaluation Results on RGBT234
Table 2: The overall performance on RGBT234. The red, blue, and green fonts indicate the first, second, and third places of each evaluation metric, respectively.
@@
M3PT 85.9 63.4 46.1fps
ECMD[42] 82.4 58.4 30.0fps
MFNet[14] 84.4 60.1 18.0fps
DMCNet[48] 83.9 59.3 2.2fps
ViPT-deep[17] 83.6 61.8 -
ViPT-shaw[17] 82.4 61.5 -
PLASSO[45] 82.7 59.2 -
APFNet[13] 82.7 57.9 1.9fps
LRMW[49] 82.5 61.6 24.6fps
MIRNet[50] 81.6 58.9 30.0fps
ADRNet[12] 80.7 57.0 25.0fps
MANet++[8] 79.5 55.9 16.1fps
M5L[7] 79.5 54.2 9.0fps
JMMAC[9] 79.0 57.4 -
HMFT[15] 78.8 56.9 25.1fps
On RGBT234, we compare our proposed method with 14 state-of-the-art methods, and the comparison results are shown in Table 2. According to Table 2, our method surpasses ViPT-deep by 2.3 and 1.6 percentage points on the MPR and MSR metrics respectively, and ranks first among all methods on all metrics and tracking speed. We also use RGBT234 toolkit and plot the MSR and MPR evaluation curves of our method and 6 open-source methods, as shown in Figure 8.
RGBT234 also provides annotations for 12 classic challenge attributes (CM, HO, FM, SV, PO, LI, NO, TC, BC, MB, DEF, LR). Therefore, we plot the per-attribute MPR evaluation curves, as shown in Figure 9. Our method shows leading performance on 10 of the 12 challenge attributes (CM, HO, FM, SV, PO, NO, TC, BC, DEF, LR), demonstrating the robustness and application potential of our method in open world scenarios.
@@ 5.4 Evaluation Results on RGBT210
Table 3: The overall performance on RGBT210. The red, blue, and green fonts indicate the first, second, and third places of each evaluation metric, respectively.
@@
M3PT 83.9 60.8 46.1fps
ViPT-deep[17] 81.3 59.2 -
ViPT-shaw[17] 80.4 59.1 -
JMMAC[9] 78.3 55.3 -
ADRNet[12] 79.4 54.8 25.0fps
DMCNet[48] 79.7 55.5 2.2fps
CAT[6] 79.2 53.3 -
MANet++[8] 77.2 53.4 25.4fps
HMFT[15] 77.7 55.7 25.1fps
We compare our method with 8 state-of-the-art methods on RGBT210, and show the comparison results in Table 3. Our method achieves PR and SR scores of 83.9 and 63.8 respectively, surpassing ViPT-deep by 2.3 and 1.6 percentage points, and demonstrating its competitiveness among all methods. We plot the SR and PR evaluation curves of our method along with six leading open-source methods, as shown in Figure 10.
@@ 5.5 Evaluation Results on VTUAV
Table 4: The overall performance on VTUAV. The red and blue fonts indicate the first and second places of each evaluation metric, respectively.
@@
M3PT 55.6 47.4 75.5 63.0 46.1fps
HMFT-LT[15] 53.6 46.1 - - 8.1fps
HMFT[15] 41.4 35.5 75.8 62.7 25.1fps
FSRPN[51] 36.6 31.4 65.3 54.4 36.8fps
mfDiMP[5] 31.5 27.2 67.3 55.4 25.8fps
DAFNet[52] 25.3 18.8 62.0 45.8 7.1fps
ADRNet[12] 23.5 17.5 62.2 46.6 10.3fps
VTUAV is the first benchmark to propose a long-term tracking task, which poses new challenges for RGB-T tracking. The VTUAV test set consists of two subsets: long-term tracking subset and short-term tracking subset. We select six RGB-T tracking methods that have been evaluated on VTUAV and compare them with our method on both subsets. Table 4 shows the results. In addition, we use the VTUAV toolkit to plot the MPR and MSR evaluation curves of our method and other open-source state-of-the-art methods on the long-term tracking subset and the short-term tracking subset, respectively, as shown in Figure 11 and Figure 12.
The evaluation results and curves show that our method performed very competitively on both subsets. It is encouraging that, despite not introducing any model online update or template update operation in the inference stage, our method still achieves state-of-the-art on the VTUAV long-term tracking test set. The second-ranked HMFT-LT method on the long-term subset, which introduces a model online update mechanism to adapt to the long-term tracking task, has an inference speed of only 8.1fps, less than one-fifth of the M3VPT inference speed. This is because our proposed method fully explores the complementary information of the dual modality and maximizes the use of upstream knowledge in the long-term tracking task, enabling it to better adapt to the more complex and open long-term tracking task.

### v2

@@ 5.3 Evaluation Results on LasHer
Table 1: The overall performance on LasHer. The best two results among each type of methods are shown in red and blue fonts, respectively.
@@
M3PT (Ours) Prompt Fine-tune 70.0 66.3 56.1
OneTrack [30] Prompt Fine-tune 67.1 63.3 54.1
UnTrack [29] Prompt Fine-tune 66.7 62.9 53.6
ViPT-deep [28] Prompt Fine-tune 64.8 61.3 52.2
ViPT-shaw [28] Prompt Fine-tune 59.0 55.7 47.6
ProTrack [27] Prompt Fine-tune 53.8 49.8 42.0
TBSI [22] Full Fine-tune 69.2 65.7 55.6
ECMD [23] Full Fine-tune 59.0 54.6 46.4
DRGCNet [18] Full Fine-tune 48.3 42.3 33.8
MFNet [16] Full Fine-tune 59.7 55.4 46.7
APFNet [15] Full Fine-tune 50.0 43.9 36.2
HMFT [17] Full Fine-tune 43.6 38.1 31.3
mfDiMP [7] Full Fine-tune 44.7 39.5 34.3
SGT++ [24] Full Fine-tune 36.5 30.6 25.1
PLASSO [58] Other 54.7 - 47.2
SiamMLAA [20] Other 53.8 - 43.1
MANet++ [10] Other 46.7 40.8 31.7
CAT [8] Other 45.0 39.5 31.4
FANet [59] Other 44.1 38.4 30.9
DAPNet [60] Other 43.1 38.3 31.4
We evaluate the performance of our proposed M3PT on LasHer, and compare it with other 19 state-of-the-art RGB-T tracking methods. To better contrast methods of the same type as well as those of differing types, we carefully divide all 20 methods into three categories: the first category consists of 6 lately excellent methods that utilize pre-trained RGB tracking models for parameter-efficient prompt fine-tuning; the second category comprises 8 cutting-edge methods that take pre-trained RGB tracking models for full-parameter fine-tuning; the third category includes 6 recent methods that do not use pre-trained upstream parameters.
[…]
LasHer provides fine annotation of 19 classic attributes, encompassing a general attribute—No Occlusion (NO)—and 18 intricate attributes: Partial Occlusion (PO), Total Occlusion (TO), Hyaline Occlusion (HO), Low Illumination (LI), High Illumination (HI), Abrupt Illumination Variation (AIV), Out-of-View (OV), Low Resolution (LR), Deformation (DEF), Background Clutter (BC), Similar Appearance (SA), Thermal Crossover (TC), Motion Blur (MB), Camera Moving (CM), Frame Lost (FL), Fast Motion (FM), Scale Variation (SV), and Aspect Ratio Change (ARC). These attributes can cover most intricate scenarios in the real world. To evaluate our proposed method in challenging scenarios, we use official toolkit to plot attribute-wise PR evaluation curves of our method and 12 other open-source methods, as shown in Figure 7. The vital conclusions could be summarized as follows: (1) In general case NO, because most sota trackers could easily track the target without any challenging disruptions, our method does not exhibit a pronounced competitive advantage against latest sota OneTrack and UnTrack. (2) In challenging cases, the advantage and robustness of our M3PT is fully demonstrated through surpassing all sota methods in 17 of the 18 challenging attributes (PO, FL, TC, SV, SA, MB, LR, LI, HI, DEF, CM, BC, ARC, AIV, TO, HO, FM), which include not only typical challenge attributes of visual tracking (TO, HO, FL, SV, PO, SA, MB, ARC, LR, CM, FM, BC, DEF), but also the extremely challenge attributes of visible modality (LI, HI, AIV) and thermal infrared modality (TC). The results proves that our method can indeed transfer upstream knowledge effectively and release great potential of multi-modal prompt tracking by uni-model exploration and scenario-adaptive fusion. (3) In the cases of OV, our tracker does not get leading performance. The result may be attributed to the amplification of location error in two branches and inter-modal fusion module, when images of both modalities all lose the target appearance. We think extra clues like motion prompts might be useful to solve this problem while maintaining our great performance on other challenging attributes.
@@ 5.4 Evaluation Results on RGBT234
Table 2: The overall performance on RGBT234. The best two results among each type of methods are shown in red and blue fonts, respectively.
@@
M3PT (Ours) Prompt Fine-tune 86.5 63.9
OneTrack [30] Prompt Fine-tune 85.7 64.2
UnTrack [29] Prompt Fine-tune 83.7 61.8
ViPT-deep [28] Prompt Fine-tune 83.6 61.8
ProTrack [27] Prompt Fine-tune 72.4 54.3
USTrack [26] Full Fine-tune 87.4 65.8
TBSI [22] Full Fine-tune 87.1 63.7
ECMD [23] Full Fine-tune 82.4 58.4
SiamIVFN [25] Full Fine-tune 81.1 63.2
DRGCNet [18] Full Fine-tune 82.5 58.1
MFNet [16] Full Fine-tune 84.4 60.1
DMCNet [61] Full Fine-tune 83.9 59.3
APFNet [15] Full Fine-tune 82.7 57.9
HMFT [17] Full Fine-tune 78.8 56.9
SiamAFTS [19] Other 89.0 60.2
PLASSO [58] Other 82.7 59.2
SiamMLAA [20] Other 79.5 58.4
SiamTDR [21] Other 77.2 55.1
LRMW [62] Other 82.5 61.6
MIRNet [63] Other 81.6 58.9
MBAFNet [64] Other 80.1 58.5
CIRNet [65] Other 81.0 54.4
ADRNet [14] Other 80.7 57.0
MANet++ [10] Other 79.5 55.9
M5L [9] Other 79.5 54.2
JMMAC [11] Other 79.0 57.4
On RGBT234, we compare our proposed method with other 24 state-of-the-art methods. Similar to evaluation on LasHer, we also divide all 26 methods into three categories, including 5 latest prompt fine-tuning methods, 9 cutting-edge full fine-tuning methods and 12 non-fine-tuning methods.
The evaluation and comparison results are shown in Table 2. (1) Compared to similar methods, our M3PT seperately ranks first and second on MPR and MSR metrics, while surpasses our baseline ViPT-deep by 2.9 and 2.1 percentage points on MPR, respectively. (2) In comparison with other types of methods, our prompt method also perform great competitiveness: on the MPR metric, M3PT notably outperforms the third-ranked full fine-tuning method and the second-ranked non-fine-tuning method, while on the MSR metric, it exceeds the performance of the second-ranked full fine-tuning method and all non-fine-tuning methods. Also, We use RGBT234 toolkit to plot the MSR and MPR evaluation curves of our method and other 6 open-source methods, as shown in Figure 8.
RGBT234 provides 12 annotated attributes, including a general attribute-NO-and 11 challenging attributes: CM, HO, FM, SV, PO, LI, TC, BC, MB, DEF, LR. We operate attribute-wise MPR evaluation, as shown in Figure 9. The conclusion could be summarized below: (1) In the general attribute, most trackers could precisely track the target and achieve a good score without challenges, therefore, our M3PT does not exhibit an obviously leading performance. (2) Our method shows leading performance on all 11 annotated challenging attributes (FM, MB, BC, TC, PO, SV, HO, DEF, LI, CM), demonstrating the generalization and application potential of our prompt-based tracking approach in open world scenarios. (3) We also observe that our method exhibits divergent performance on the same attributes across RGBT234 and LasHeR. For instance, concerning the FM and BC attributes in RGBT234, our method’s improvement relative to the baseline is not as pronounced as observed in LasHeR. This discrepancy may be ascribed to divergent training sets employed for model evaluation across the two benchmarks. Additionally, observing that our baseline and other sota methods exhibit similar phenomenon, we think variations in data distribution between the two benchmarks, such as the presence of different object categories and the cross-distribution of distinct attributes within the same sequences, might also contribute to the inconsistency.
@@ 5.5 Evaluation Results on RGBT210
Table 3: The overall performance on RGBT210. The best two results among each type of methods are shown in red and blue fonts, respectively.
@@
M3PT (Ours) Prompt Fine-tune 85.0 61.3
UnTrack [29] Prompt Fine-tune 81.5 59.0
ViPT-deep [28] Prompt Fine-tune 81.3 59.2
ProTrack [27] Prompt Fine-tune 69.7 52.0
TBSI [22] Full Fine-tune 85.3 62.5
DMCNet [61] Full Fine-tune 79.7 55.5
HMFT [17] Full Fine-tune 77.7 55.7
SiamMLAA [20] Other 77.9 56.7
CIRNet [65] Other 78.9 52.3
ADRNet [14] Other 79.4 54.8
JMMAC [11] Other 78.3 55.3
MANet++ [10] Other 77.2 53.4
CAT [8] Other 79.2 53.3
On RGBT210, We compare our method with other 12 state-of-the-art methods, including 3 prompt fine-tuning methods, 3 full fine-tuning methods and 6 non-fine-tuning methods. As shown in Table 3, our method surpasses the baseline ViPT-deep by 3.7 and 2.1 percentage points, and ranks first among prompt methods on all metrics. Furthermore, it demonstrates a PR score that closely approaches the best full fine-tuning method. The SR and PR evaluation curves are shown in Figure 10.
5.6 Evaluation Results on GTOT
Table 4: The overall performance on GTOT. The best result among each type of methods are shown in red fonts.
[…]
M3PT (Ours) Prompt Fine-tune 92.7 76.9
ViPT-deep [28] Prompt Fine-tune 91.8 76.1
ProTrack [27] Prompt Fine-tune 80.8 68.9
USTrack [26] Full Fine-tune 93.4 78.3
SiamIVFN [25] Full Fine-tune 91.5 79.3
ECMD [23] Full Fine-tune 90.7 73.5
DRGCNet [18] Full Fine-tune 90.5 73.5
DMCNet [61] Full Fine-tune 90.9 73.3
HMFT [17] Full Fine-tune 91.2 74.9
APFNet [15] Full Fine-tune 90.5 73.7
SiamMLAA [20] Other 91.3 75.1
SiamTDR [21] Other 88.5 71.4
SiamAFTS [19] Other 84.9 77.7
MIRNet [63] Other 90.9 74.4
CIRNet [65] Other 90.1 72.8
ADRNet [14] Other 90.4 73.9
JMMAC [11] Other 90.2 73.2
MANet++ [10] Other 88.2 70.7
CAT [8] Other 88.9 71.7
On GTOT, We compare our method with other 17 state-of-the-art methods, including 2 prompt fine-tuning methods, 6 full fine-tuning methods and 9 other methods. Our M3PT reaches 92.7 PR percentage points and 76.9 SR percentage points. When reaching the best result among prompt fine-tuning methods, M3PT also outperforms the best non-fine-tuning method and the second-best full fine-tuning in terms of PR, as shown in Table 4. We also draw the evaluating metric plots of 10 open-source methods, as shown in Figure 11.
@@ 5.7 Evaluation Results on VTUAV
Table 5: The overall performance on VTUAV. The best result among each type of methods are shown in red fonts.
Type Long-term Short-term
[…]
M3PT (Ours) Prompt Fine-tune 55.6 47.4 75.5 63.0
ViPT-deep [28] Prompt Fine-tune 54.1 46.5 74.6 62.3
ProTrack [27] Prompt Fine-tune 46.7 41.7 72.5 62.9
HMFT-LT [17] Full Fine-tune 53.6 46.1 - -
HMFT [17] Full Fine-tune 41.4 35.5 75.8 62.7
mfDiMP [7] Full Fine-tune 31.5 27.2 67.3 55.4
FSRPN [66] Other 36.6 31.4 65.3 54.4
DAFNet [67] Other 25.3 18.8 62.0 45.8
ADRNet [14] Other 23.5 17.5 62.2 46.6
VTUAV is the first benchmark to propose a long-term tracking subtask. To evaluate our method on this benchmark, we select 8 methods, including 2 prompt fine-tuning methods, 3 full fine-tuning methods and 3 none-fine-tuning methods, and compare them with our M3PT on long-term and short-term subsets, as shown in Table 5. In addition, we plot MPR and MSR evaluation curves on two subsets by official tookit, respectively, as shown in Figure 12 and Figure 13.
It is encouraging that, despite not introducing any model online update or template update operation in the inference stage, our method still achieves state-of-the-art among all the methods on long-term subset, while the best full fine-tuning method, HMFT-LT, introduces an online parameter-update mechanism to adapt to the long-term tracking task. This is because our proposed method could fully explore the complementary information of the dual modality and maximizes the use of excellent upstream knowledge, ensuring robust performance when coping with long-term challenge.
5.8 Evaluation Results on VOT2019RGB-TIR
Table 6: The overall performance on VOT2019RGB-TIR. The best result among each type of methods are shown in red fonts.
[…]
M3PT (Ours) Prompt Fine-tune 67.2 81.3 41.6
ViPT-deep [28] Prompt Fine-tune 66.1 73.1 34.8
ProTrack [27] Prompt Fine-tune 63.7 71.0 34.2
DMCNet [61] Full Fine-tune 60.1 70.9 38.0
SiamCDA [12] Full Fine-tune 68.2 75.7 42.4
JMMAC [11] Other 66.0 82.4 49.8
TFNet [13] Other 46.2 59.4 28.8
ADRNet [14] Other 62.2 75.7 39.6
MANet++ [10] Other 50.9 53.8 27.2
FSRPN [66] Other 63.6 70.7 35.5
Different with other mainstream benchmarks, VOT2019RGB-TIR introduce EAO, Robustness and Accuracy metrics to evaluate the performance of RGB-T Trackers. On VOT2019RGB-TIR, we compare our method with other 9 state-of-the-art methods, including 2 prompt fine-tuning methods, 2 full fine-tuning methods and 5 other methods.
[…]
5.9 Fine-tuned Parameters Comparison
Table 7: The parameters comparison of different fine-tuning methods. Params represents the number of whole parameters of the model, and Tuned Params represents the number of unfrozen parameters involved in fine-tuning.
Method Type Total Params Tuned Params LasHer
[…]
RGB Tracker Foundation model 92.12M - 51.5 41.2
@@ 5.9 Fine-tuned Parameters Comparison
[…]
To verify the parameter-efficiency of our prompt fine-tuning method, we compare it with other recent prompt tine-tuning methods and state-of-the-art full-parameter fine-tuning methods again in terms of parameters and performance. To ensure a fair comparison, all fine-tuning method are built upon the same RGB tracking method [40].
[…]
5.10 Ablation Study
Table 8: The ablation experiment results. In the table, the four abbreviations represent the four prompting strategies. UIEPS: Uni-modal and Inter-modal Exploration Prompt Strategy, MFPS: Middle Fusion Prompt Strategy, FEPS: Fusion-modal Enhancing Prompt Strategy, MSPS: Modality-aware and Stage-aware Prompt Strategy.
UIEPS MFPS FEPS MSPS RGBT234
[…]
69.5 52.1
$\surd$ 74.9 56.5
$\surd$ 80.2 59.8
$\surd$ $\surd$ 83.4 61.7
$\surd$ 76.7 56.8
$\surd$ 72.3 54.2
$\surd$ $\surd$ $\surd$ $\surd$ 85.1 63.0
$\surd$ $\surd$ $\surd$ $\surd$ $\surd$ 86.5 63.9
To verify the effectiveness of our proposed four prompt strategies, we conduct extensive ablation studies and analyses on RGBT234. In the experiments, we build a baseline tracker which is only composed of upstream parameters, with MFP replaced by element-wise addition operation of dual-modal features. Based on the baseline, we explore the effectiveness of different prompt strategies. The ablation results are shown in Table 8.
Uni-modal and Inter-modal Exploration Prompt Strategy: We investigate the impact of UEP, IP and complete Uni-modal and Inter-modal Exploration Prompt Strategy on our M3PT. As shown in Table 8, UEP brings a guidance for the upstream backbone to extract comprehensive uni-modal features while IP leads to effective inter-modal communication by channel-level prompts. Compared with base tracker, the collaboration of UEP and IP improves the MPR and MSR of base tracker on RGBT234 by 13.9 and 9.6 percentage points, respectively, which proves that this strategy significantly guides the backbone to fully explore modality-shared features and modality-independent features.
Middle Fusion Prompt Strategy: We only utilize MFP on the base tracker, to verify the effectiveness of this strategy. According to Table 8, the model only with Middle Fusion Prompt Strategy surpasses base tracker by 7.2 and 4.7 percentage points on the MPR and MSR metrics on RGBT234, respectively. This indicates that this prompt strategy can achieve adaptive selection of discriminative features, providing reliable fusion prompts for the second-stage backbone.
Fusion-modal Enhancing Prompt Strategy: We validate Fusion-modal Enhancing Prompting Strategy. Table 8 shows that this strategy boosts the MPR and MSR metrics of the original tracker on RGBT234 by 2.8 and 2.1 percentage points, respectively. This confirms that the strategy better adapts the second-stage backbone to fusion-modal modeling and feature enhancement.
Modality-aware and Stage-aware Prompt Strategy: To evaluate the effectiveness and necessity of the Modality-aware and Stage-aware Prompt strategy, we build a tracker that contains the other three prompt strategies and a tracker that contains all four prompt strategies. As shown in Table 8, although the collaboration of the three prompt strategies for different purposes has already achieved excellent performance on RGBT234, this strategy still improves MPR and MSR metrics by 1.4 and 0.9 points, respectively. With only a small amount of parameters, the strategy stores low-level fixed patterns within the modality and enables the foundation model to better identify current modality and sub-task.
5.11 Variable Analysis
Table 9: Fusion location in our proposed middle fusion tracking framework.
[…]
1 62.1 47.3 94.7fps
6 67.8 51.4 71.3fps
11 73.5 55.2 59.1fps
Variable in Middle Fusion Framework. We explore the impact of different fusion location value on speed and performance of tracker with our middle fusion meta-framework. First, we remove all prompt modules like the base tracker in ablation study, to prevent the influence of other components. From Table 9, the tracker operates at a fast speed with marginally inferior performance when the fusion module is inserted after the 1st layer. Conversely, the inference speed yields to enhanced performance when the fusion module is inserted after the 11th layer. When the fusion location is set to 6, the tracker’s performance and speed have achieved a harmonious balance. These results show our Middle fusion Framework’s flexibility, indicating that we could adjust the variable fusion location to strike an relatively ideal balance between effectiveness and efficiency, thereby meeting the diverse requirements of various applying scenarios.
Table 10: First-stage block number in our proposed M3PT.
[…]
3 85.1 62.9 44.6fps
6 85.5 63.3 39.9fps
10 86.5 63.9 34.1fps
First-Stage Block Number of M3PT. Considering our method sets an example for application of the proposed middle fusion meta-framework, we also analysis the impact of the first-stage block number on our M3PT. From Table 10, adding the number of the first-stage blocks augments the extraction of uni-modal information and fosters a more effective complementary interplay between the two modalities, thereby refining the foundational model’s adaptation to the downstream task, albeit at the expense of efficiency. The findings verify the flexibility of our fusion framework again.